Use the given degree of confidence and sample data to construct a confidence interval for the population proportion

p.

nequals=​195,

xequals=​162;

​95% confidence

A.

0.777less than<pless than<0.884

B.

0.789less than<pless than<0.873

C.

0.788less than<pless than<0.873

D.

0.778less than<pless than<0.883
Answer #1

Solution :

Given that,

n = 195

x = 162

Point estimate = sample proportion = \hat p = x / n = 162 / 195 = 0.83

1 - \hat p = 1 - 0.83 = 0.17

At 95% confidence level the z is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

\alpha / 2 = 0.05/ 2 = 0.025

Z\alpha/2 = Z 0.025 = 1.645

Margin of error = E = Z\alpha / 2 * \sqrt ((\hat p * (1 - \hat p )) / n)

= 1.96 * (\sqrt((0.17*0.83)/ 195)

= 0.053

A 95% confidence interval for population proportion p is ,

\hat p - E < p < \hat p + E

0.831 - .053 < p < 0.831+ 0.53

0.778 < p < 0.883

Option (D) is correct